OpenStudy (anonymous):

In triangle ABC, the measure of B is 6 degrees more than three times the measure of angle A. The measure of angle C is 39 degrees more than the measure of angle A. Find the measure of each angle.

OpenStudy (anonymous):

Someone help please! :)

OpenStudy (anonymous):

let angle A be 'x' degrees. then angle B would be '3x+6' degrees and angle C would be 'x+39'. sum up these angles and equate them to 180. then solve for x

OpenStudy (ny,ny):

you got it?

OpenStudy (ny,ny):

or you want more explanation?

OpenStudy (anonymous):

Yeah, I need the answer for A, B, and C. Please!

OpenStudy (ny,ny):

so 4x + 45 = 180 after calculations you get: 4x = 135 More calculations... x = 33.75

OpenStudy (ny,ny):

x= a

OpenStudy (anonymous):

Huh? SHouldn't it be a number?

OpenStudy (ny,ny):

33.75 = a 107.25 = b 72.75 = c

OpenStudy (ny,ny):

that kinda number?

OpenStudy (anonymous):

Yeah, but rounded I think?

OpenStudy (ny,ny):

rounded?

OpenStudy (anonymous):

I just tried the first answer and it was wrong. :(

OpenStudy (ny,ny):

my answer?

OpenStudy (anonymous):

Yeah.

OpenStudy (ny,ny):

oh

OpenStudy (ny,ny):

i think i messed up one sec

OpenStudy (anonymous):

Oh ok!

OpenStudy (ny,ny):

my bad it was 5x + 45 = 180

OpenStudy (ny,ny):

im sorry

OpenStudy (ny,ny):

a = 27 b= 86 (i think this is all in my head) c = 46

OpenStudy (ny,ny):

wait im wrong

OpenStudy (ny,ny):

im so dumb!!!!

OpenStudy (ny,ny):

a = 27 b= 87 c = 66

OpenStudy (ny,ny):

now im sure

OpenStudy (ny,ny):

27 + 66 + 87 = 180 YAY IM RIGHT!!!!!!

OpenStudy (anonymous):

Oh wow. You are the man! Thank you so much!

OpenStudy (ny,ny):

No. It was nothing. Your a pretty girl, I just had to help you out.

OpenStudy (ny,ny):

Bye

OpenStudy (anonymous):

Awe thanx!

OpenStudy (ny,ny):

sorry for my failures